# Quantum Pontus-Mpemba Effects in Real and Imaginary-time Dynamics

### v1

@@ IV Discussions and Conclusion
In this study, we introduce and investigate the concept of quantum Pontus-Mpemba effects, a counterintuitive phenomenon where a quantum system relaxes faster through a two-step evolution protocol compared to a direct quench. By employing both real-time and imaginary-time dynamics, we demonstrate that QPME emerges in systems with $U(1)$ symmetry, particularly when initialized in tilted ferromagnetic states at small angles. In real-time, the asymmetric evolution of the two-step protocol increases the thermal nature of the initial state, resulting in significantly accelerating thermalization. For imaginary-time evolution, the two-step protocol also accelerates convergence to the ground state for small tilt angles. We attribute this speedup to the broader charge distribution generated by the initial asymmetric evolution. Conversely, QPME is suppressed for large tilt angles or for antiferromagnetic initial states. Notably, our finite-size scaling analysis shows that the QPME persists across different system sizes, indicating its robustness in the thermodynamic limit.
@@
While of fundamental interest, the true strength of the QPME protocol lies in its practical applicability. Whereas QME compares the dynamics evolving from different initial states to the same target state, the QPME protocol relies on just a single initial state. This distinction is operationally significant: in many experimental settings, the central challenge is not to compare different starting conditions, but to identify the most efficient dynamical path from a given initial state toward a final target state. The QPME protocol directly addresses this need by leveraging transient symmetry breaking to accelerate the relaxation process. Furthermore, the principles of QPME in imaginary time dynamics suggest a powerful new strategy for enhancing the performance of numerical algorithms. For tensor network methods like TEBD, a shorter evolution time translates to fewer computational steps, leading to faster ground state preparation. This speed-up is also particularly impactful for projection quantum Monte Carlo (PQMC) simulations. In many interesting systems, PQMC is plagued by the infamous sign problem, where computational complexity grows exponentially with the projection time. By achieving convergence in a shorter imaginary time, our protocol can substantially mitigate the sign problem [81, 82, 83, 84, 85, 86, 87, 88, 89], reducing the computational resources required and potentially extending the reach of PQMC to previously intractable problems.
@@
Hence, our findings not only elucidate a fundamental non-equilibrium phenomenon but also provide a valuable strategy for accelerating quantum simulations and state preparation. Future work could extend these results to other symmetries, open quantum systems, or more complex interaction geometries.
@@ ACKNOWLEDGMENTS
The work is supported by the Ministry of Science and Technology (Grant No. 2022YFA1403900), the Strategic Priority Research Program of the Chinese Academy of Sciences (Grant Nos. XDB28000000, XDB33000000), and the New Cornerstone Investigator Program. HY is supported by the International Young Scientist Fellowship of the Institute of Physics, Chinese Academy of Sciences (No.202407). SXZ acknowledges the support from the Innovation Program for Quantum Science and Technology (2024ZD0301700) and the start-up grant at IOP-CAS.

### v2

@@ V Discussions and Conclusion
In this study, we introduce and investigate the concept of quantum Pontus-Mpemba effects, a counterintuitive phenomenon where a quantum system relaxes faster through a two-step evolution protocol compared to a direct quench. By employing both real-time and imaginary-time dynamics, we demonstrate that QPME emerges in systems with $U(1)$ symmetry, particularly when initialized in tilted ferromagnetic states at small angles. In real-time, the asymmetric evolution of the two-step protocol increases the thermal nature of the initial state, resulting in significantly accelerating thermalization. For imaginary-time evolution, the two-step protocol also accelerates convergence to the ground state for small tilt angles. We attribute this speedup to the broader charge distribution generated by the initial asymmetric evolution. Conversely, QPME is suppressed for large tilt angles or for antiferromagnetic initial states. Our variational analysis in Sec. IV further confirms that this acceleration is not limited to specific heuristically chosen Hamiltonians; rather, an optimized symmetry-breaking path exists that enhances relaxation efficiency. Notably, our results for system sizes up to $L=48$ show consistent QPME behavior, which provides supportive evidence of its potential robustness in the thermodynamic limit.
@@
While of fundamental interest, the true strength of the QPME protocol lies in its practical applicability. Whereas QME compares the dynamics evolving from different initial states to the same target state, the QPME protocol relies on just a single initial state. This distinction is operationally significant: in many experimental settings, the central challenge is not to compare different starting conditions, but to identify an efficient dynamical path from a given initial state toward a final target state. The QPME protocol directly addresses this need by leveraging transient symmetry breaking to accelerate the relaxation process. Furthermore, the principles of QPME in imaginary time dynamics suggest a powerful new strategy for enhancing the performance of numerical algorithms. For tensor network methods like TEBD, a shorter evolution time translates to fewer computational steps, leading to faster ground state preparation. This speed-up is also particularly impactful for projection quantum Monte Carlo (PQMC) simulations. In many interesting systems, PQMC is plagued by the infamous sign problem, where computational complexity grows exponentially with the projection time. By achieving convergence in a shorter imaginary time, our protocol can substantially mitigate the sign problem [43, 60, 97, 79, 10, 47, 55, 56, 85], reducing the computational resources required and potentially extending the reach of PQMC to previously intractable problems.
@@
Hence, our findings not only elucidate an intriguing non-equilibrium phenomenon but also provide a valuable strategy for accelerating quantum simulations and state preparation. Future work could extend these results to other symmetries, open quantum systems, or more complex interaction geometries.
@@ ACKNOWLEDGMENTS
The work is supported by the Ministry of Science and Technology (Grant No. 2022YFA1403900), the Strategic Priority Research Program of the Chinese Academy of Sciences (Grant Nos. XDB28000000, XDB33000000), and the New Cornerstone Investigator Program. HY is supported by the International Young Scientist Fellowship of the Institute of Physics, Chinese Academy of Sciences (No.202407). SXZ acknowledges the support from the Innovation Program for Quantum Science and Technology (2024ZD0301700) and the National Natural Science Foundation of China (No. 12574546).